OpenStudy (anonymous):

The 2010 census reported about 308.7 million people in the United States. The 2000 census reported a population of about 281.4 million. By about what percent did the population change? Did the population grow or shrink? Hint: To find the answer, find the difference of the populations.

OpenStudy (amistre64):

whats the difference in the pops?

OpenStudy (amistre64):

\[\%change=\frac{end-start}{start}\]

OpenStudy (anonymous):

308.7 million - 281.4 million = 27,300,000 I believe. Sooo, I would divide 27,300,000 / 281,400,000 at that point?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Alright! That equals 0.09701492537, which should be rounded to 0.097 I think. Is this the number I turn into a percentage?

OpenStudy (amistre64):

308.7 million - 281.4 million -------------------------- 281.4 million (308.7 -281.4)/281.4 = .097... and yes, thats what you would want to convert to percentage

OpenStudy (anonymous):

Okay, to do that, I multiply by 100, right? .097 * 100= 9.7%! Which also means that the population grew! And that's my answer?

OpenStudy (amistre64):

yep, thats it :)

OpenStudy (anonymous):

Yay!! Thank you so much, you were very helpful. I'm understanding this sort of problem a lot better now!